OpenStudy (anonymous):

S=2兀r2+2兀rh solve for h

OpenStudy (goformit100):

What the given equation is about ?

OpenStudy (nincompoop):

how about you solve it for us and show us your solution, we will check if you are wrong along the way

OpenStudy (anonymous):

I dont know how..i have to solve for h

OpenStudy (nincompoop):

how could you not know how?

OpenStudy (anonymous):

I mean that why i need help

OpenStudy (nincompoop):

do you have any idea what to do at all, or you want us to replace your teacher for you?

OpenStudy (anonymous):

Im trying to help my friend with her algebra 2 homework

OpenStudy (anonymous):

Wow your nice.

OpenStudy (nincompoop):

A = B + C to solve for B, we isolate it by subtracting C on both sides of the equation giving us A-C = B + C - C performing the operation, we end up with A - C = B the same concept occurs in your problem. first, identify the operations, obtain their opposite in order to isolate that which you are looking for.

OpenStudy (anonymous):

-.-

OpenStudy (anonymous):

Im still lost

OpenStudy (nincompoop):

the opposite of multiplication is division and addition is subtraction. and vice versa

OpenStudy (nincompoop):

if you are lost then, you are missing a lot of fundamental knowledge prior to arriving at the current level of math you are taking.

OpenStudy (anonymous):

Thanks, i mean not really im a freshman taking algebra 2 honors and its a hard class

OpenStudy (nincompoop):

let us put those letters A, B and C into actual numerical values. 25 = 18 + 7 if we want to look for 18 we isolate 18 by subtracting 7 on both sides of the equations 25 - 7 = 18 - 7 performing the operation we will obtain 25 - 7 = 18 continue 18 = 18 this is the simplest way of explaining the concept you are missing. instead of letters they became actual values and it could be more letters and constant values - values that do not change like pi, which we normally refer to simplest value as 3.14

OpenStudy (nincompoop):

wrong answer. you are to inquire that which you do not understand, but you have to be specific

OpenStudy (nincompoop):

S=2兀r^2+2兀rh you want to look for the value of h the h is stuck together with 2兀rh but for the time being we can imagine 2兀r^2+2兀rh like B + C we want to isolate B first, so you need to subtract B which is 2兀r^2 can you at least perform that task and show me what you've come up with?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

S-2兀r2= S= 2兀rh

OpenStudy (anonymous):

Wrong

OpenStudy (anonymous):

S-2兀r2=2兀rh

OpenStudy (anonymous):

Do i divide 2兀r?

OpenStudy (nincompoop):

I think you got the idea now let us break down 2兀rh to isolate h remember that 2兀rh is all multiplication and we need its opposite, which is division would it make sense if we divide 2兀r on both sides of the equation

OpenStudy (nincompoop):

see I think you're a smart kid! how about you do what you thought you should

OpenStudy (anonymous):

I just gave low self esteem

OpenStudy (nincompoop):

what do we have next?

OpenStudy (anonymous):

S-2兀r2/2兀r=h

OpenStudy (anonymous):

What do you mean by what do we have next?

OpenStudy (nincompoop):

yes I mean what you have now is what I was asking

OpenStudy (anonymous):

So thats correct?

OpenStudy (anonymous):

I have another problem

OpenStudy (nincompoop):